WELL I KNOW HOW TO GET THE VERTEX OF A QUADRATICS EQUATION BUT I DONT GET HOW TO GET OR FIND THE ROOT FOR EXAMPLE FINDIN THE ROOT OF THIS PROBLEM:
Y=(x-2)(x+6)

Answer by stanbon(48558)

The "root" is an "x" value that has a corresponding "y" value of zero.
So let (x-2)(x+6)=0
Then either x-2=0 or x+6=0
which means x=2 or x=-6
They are the "root" values of your equation.
Cheers,
stan H.